MICRO Inc. has net income of $74 million. It has identified positive NPV projects that require $66.6 million in funding. The company's target debt/equity ratio is 0.5.

Part 1. What is the company's target capital structure weight for equity (E/V)?

Part 2. If the company follows a residual dividend policy, how much will it need to borrow (in $ million)?

Part 3. If the company follows a residual dividend policy, how much will it pay out in dividends (in $ million)?
Answer #1

Given

Net Income = $74million
NPV (Capital Expenditure required) = 66.6million

Debt/Equity = .5

1.

Weight of Equity

For weight of equity we have
Debt/ Equity = 1/2 - Equation 1
Debt + Equity will be 1+2 = 3 - Equation 2

From Equation 1 and 2
Weight of Equity = 2/3 = 66.667% =.667

2.

Amount to be borrowed

For the amount of debt to be borrowed we have
Debt/ Equity = 1/2 - Equation 1
Debt + Equity will be 1+2 = 3 - Equation 2

From Equation 1 and 2

Weight of Debt = 1/3 = 33.333% = .333

Capital Expenditure = 66.6 million

Debt = Weight of Debt * Capital Expenditure

=.3333*66.6

Amount of Debt to be = 22.2 million

3.

The company follows a residual dividend policy,
According to residual dividend policy
Dividend = Net Income - [Weight of Equity*Capital Expenditure]
= 74-[66.667%*66.6]

= 29.6 million Paid out as dividend